OpenStudy (anonymous):

@tkhunny If you pour a cup of coffee that is 200 F, and set it on desk in a room that is 68 F, and 10 minutes later it is 145 F, what temperature will it be 15 minutes after you originally poured it?

OpenStudy (tkhunny):

You will need a model of some sort. Are we using Newton's Law of Cooling or some simpler function for our cooling process?

OpenStudy (anonymous):

In my class, we talked about Newton's Law of Cooling. The formula is T = A + Se^(-kt) right?

OpenStudy (tkhunny):

There you go. Please assign values to the known parameters. T = ?? A = ?? S = ?? e = ?? k = ?? t = ??

OpenStudy (anonymous):

T= Would this be the temperature after 15 minutes? A= 68 F S= 200 F k= constant? t= 15 minutes

OpenStudy (tkhunny):

A is the Ambient Temperature. This is 68ºF S is the Initial Difference in Temperature. This is 200ºF - 68ºF = 132ºF These are the two important parameters to know with surety. Our task is to find 'k'. This is normally done with a hint. It had better be given in teh problem statement. " and 10 minutes later it is 145 F, " THERE it is!! Knowing that T(10 min) = 145ºF = 68ºF + 132ºF * e^(-k*(10 min)), solve for k and you'll be nearly done.

OpenStudy (anonymous):

Ah! Most of it makes sense, but solving for k is confusing to me. How do I solve for it?

OpenStudy (tkhunny):

You should have enough logarithms in your background. No?

OpenStudy (anonymous):

I'm really not good at math, so no, I'm not good with any of this.

OpenStudy (tkhunny):

Then you cannot solve this problem. There should have been prerequisites to this course. Those prerequisites would have included simple problem solving and logarithms. Solve for P 112 = 61 + 97P

OpenStudy (anonymous):

P= 51/97 or P=52 You wanted me to solve that right? Or was it an example?

OpenStudy (tkhunny):

I did intend for you to do that. Hopefully, you meant P = 0.52 Okay, now go solve your problem. It requires EXACTLY the same process.

OpenStudy (anonymous):

So, T(15)= 126.9 F? And yes, I did mean 0.52 whoops.